OpenStudy (anonymous):

Which test of convergence should I use? sum((5-2sqrt(n))/(n^3),x,1,infinity)?

OpenStudy (anonymous):

Better format \[\sum_{1}^{\infty}(5-2\sqrt{n})/n^{3}\]

OpenStudy (anonymous):

If you break the series into \[\sum_{1}^{\infty}5/n^3+\sum_{1}^{\infty}2/n^2.5\] You'll find that both these series converge by the p-series test. Since both of these converge, the sum of the two is convergent.

OpenStudy (anonymous):

Sorry, there's a minus, not a plus... but the result is the same.

OpenStudy (anonymous):

Oh wow. That makes a lot of sense. haha I was trying to use integral rule, which wasn't going too well. Thanks!